# Kernel Search approach to solve the Minimum Spanning Tree Problem with conflicting edge pairs00footnotetext: Email addresses:

### v1

@@ 5 Conclusion
In this paper, we introduced a novel solution approach to the Minimum Spanning Tree with Conflict problem, leveraging the Kernel Search method [1]. This algorithm focuses on obtaining a feasible solution, ideally of high quality, by strategically selecting a small set of promising variables. This subset, known as the kernel set, is initially constructed through the solution of the LP relaxation of the original ILP. Throughout the algorithm, the kernel set is dynamically enriched by identifying new promising variables through the solution of ILP subproblems of increasing size. A key innovation in our methodology lies in the incorporation of independent sets of proper subgraphs within the KS algorithm. This novel approach draws inspiration from the observation that the edges of any feasible MSTC solution correspond to an independent set in the conflict graph.
Through comparative analysis with existing heuristics, we demonstrated the efficacy of our method in producing competitive results. It is evident that our tailored Kernel Search approach offers a promising tool for tackling the MSTC problem and potentially other conflict-based optimization challenges.
Acknowledgement
The research of F. Carrabs and M. Cerulli was partially supported by the project “SEcurity and RIghts in the CyberSpace” SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU. This support is gratefully acknowledged.
Appendix
In this section, we show the detailed results of the comparison among the HDA+, ClassicKS, and KS on the CCPR instances proposed in [6]. As already noted for the other tables, whenever the B&C method finds the optimal value, we report such value with the “*” symbol in the UB_best column. Whenever no feasible solution is known in the literature, but we do find one, a “-” is reported in the UB_best column. In the gap columns, if the considered method does not find a feasible solution, we report a $100$. Instead, if the considered method does find a feasible solution, but there is a “-” in the UB_best column, we report a $-100$. Also in this case like for the previous tables, we report the results only on the instances where at least one of the methods (including ours) finds a feasible solution. As a consequence, the instances with ID 151–155, 166–170, 181, 182, 184, 185, 196–200, 211–215, 226–230 are not present in this table. The headings of Table 3 are the same as in Tables 1-2. Also in this table, we report the “Average values” and the “%best” rows at the end of the table. In both the computation of the “Average values” and the “%best” we exclude the instances in which either none of the methods in literature find a feasible solution (a “-” is reported in the UB_best column) or the KS method does not find it (the gap is 100 in its column). This happens only for four instances: 183, 191, 192, and 193.
Table 3: Comparison of the results on the CCPR instances proposed in [6], for which at least a feasible solution is found.
@@ Acknowledgement
 ID $n$ $m$ $|C|$   UB_best   time gap  time gap  time gap
 51 25 60 18   $*{}^{*}$   0.03 0.00  0.02 0.00  0.02 0.00
 52 25 60 18   $*{}^{*}$   0.03 0.00  0.00 0.00  0.00 0.00
 53 25 60 18   $*{}^{*}$   0.03 0.00  4.15 1.98  4.47 0.00
 54 25 60 18   $*{}^{*}$   0.03 0.00  0.00 0.00  0.00 0.00
 55 25 60 18   $*{}^{*}$   0.03 0.00  0.00 0.00  0.00 0.00
 56 25 60 71   $*{}^{*}$   0.04 0.00  0.42 0.00  0.52 0.00
 57 25 60 71   $*{}^{*}$   0.04 0.00  0.62 0.26  1.19 0.26
 58 25 60 71   $*{}^{*}$   0.04 0.00  0.02 0.00  0.02 0.00
 59 25 60 71   $*{}^{*}$   0.04 0.00  0.02 0.00  0.02 0.00
 60 25 60 71   $*{}^{*}$   0.04 0.00  0.00 0.00  0.00 0.00
 61 25 60 124   $*{}^{*}$   0.04 0.00  0.00 0.00  0.00 0.00
 62 25 60 124   $*{}^{*}$   0.04 0.00  0.00 0.00  0.00 0.00
 63 25 60 124   $*{}^{*}$   0.06 3.49  0.20 0.00  0.53 0.00
 64 25 60 124   $*{}^{*}$   0.05 0.00  0.26 1.25  0.68 0.00
 65 25 60 124   $*{}^{*}$   0.08 0.24  0.12 0.00  0.40 0.00
 66 25 90 41   $*{}^{*}$   0.03 0.00  2.80 0.00  2.84 0.00
 67 25 90 41   $*{}^{*}$   0.04 0.00  0.00 0.00  0.00 0.00
 68 25 90 41   $*{}^{*}$   0.03 0.00  0.69 0.00  0.84 0.00
 69 25 90 41   $*{}^{*}$   0.03 0.00  5.08 4.38  1.89 0.00
 70 25 90 41   $*{}^{*}$   0.03 0.00  12.87 0.00  21.12 0.00
 71 25 90 161   $*{}^{*}$   0.06 0.00  0.05 0.00  0.05 0.00
 72 25 90 161   $*{}^{*}$   0.07 0.00  0.00 0.00  0.00 0.00
 73 25 90 161   $*{}^{*}$   0.06 0.00  0.02 0.00  0.02 0.00
 74 25 90 161   $*{}^{*}$   0.07 0.61  0.10 0.30  0.29 0.30
 75 25 90 161   $*{}^{*}$   0.07 0.31  0.59 1.53  1.74 0.00
 76 25 90 281   $*{}^{*}$   0.08 0.00  0.25 0.00  0.41 0.00
 77 25 90 281   $*{}^{*}$   0.10 0.00  0.91 0.00  1.22 0.00
 78 25 90 281   $*{}^{*}$   0.09 0.00  0.17 0.30  0.38 0.00
 79 25 90 281   $*{}^{*}$   0.12 2.87  0.22 0.00  0.65 0.00
 80 25 90 281   $*{}^{*}$   0.11 0.56  0.16 0.00  0.43 0.00
 81 25 120 72   $*{}^{*}$   0.04 0.00  0.00 0.00  0.00 0.00
 82 25 120 72   $*{}^{*}$   0.04 0.00  0.97 0.00  7.51 0.00
 83 25 120 72   $*{}^{*}$   0.05 0.00  0.00 0.00  0.00 0.00
 84 25 120 72   $*{}^{*}$   0.04 0.00  2.56 0.00  1.18 0.00
 85 25 120 72   $*{}^{*}$   0.04 0.00  0.00 0.00  0.00 0.00
 86 25 120 286   $*{}^{*}$   0.10 0.00  0.14 0.31  0.38 0.00
 87 25 120 286   $*{}^{*}$   0.11 0.00  0.51 0.00  0.53 0.00
 88 25 120 286   $*{}^{*}$   0.08 0.00  0.00 0.00  0.00 0.00
 89 25 120 286   $*{}^{*}$   0.11 0.32  0.07 0.00  0.26 0.00
 90 25 120 286   $*{}^{*}$   0.08 0.00  0.02 0.00  0.02 0.00
 91 25 120 500   $*{}^{*}$   0.17 3.65  0.71 1.82  4.49 0.61
 92 25 120 500   $*{}^{*}$   0.16 2.36  2.83 0.00  1.17 0.00
 93 25 120 500   $*{}^{*}$   0.19 4.08  1.50 0.82  0.68 0.00
 94 25 120 500   $*{}^{*}$   0.14 0.96  0.14 0.00  0.38 0.00
 95 25 120 500   $*{}^{*}$   0.16 1.25  0.19 0.62  0.58 0.62
 96 50 245 299   $*{}^{*}$   0.24 0.00  4.27 0.00  211.21 0.00
 97 50 245 299   $*{}^{*}$   0.24 0.00  0.29 0.00  1.74 0.00
 98 50 245 299   $*{}^{*}$   0.22 0.00  0.01 0.00  0.01 0.00
 99 50 245 299   $*{}^{*}$   0.26 0.00  0.35 0.16  1.57 0.00
 100 50 245 299   $*{}^{*}$   0.28 0.00  21.59 0.00  5.83 0.00
 101 50 245 1196   $*{}^{*}$   0.62 2.95  10.90 0.00  3.96 0.00
 102 50 245 1196   $*{}^{*}$   0.64 5.87  86.60 3.08  4.61 0.00
 103 50 245 1196   $*{}^{*}$   0.64 2.26  17.80 0.71  10.42 0.00
 104 50 245 1196   $*{}^{*}$   0.55 2.66  11.05 0.16  9.39 0.00
 105 50 245 1196   $*{}^{*}$   0.65 9.84  23.49 0.59  37.74 0.29
 106 50 245 2093   833   1.00 100  1102.05 1.20  456.01 -1.44
 107 50 245 2093   835   1.02 100  3685.50 100  452.01 0.00
 108 50 245 2093   840   0.91 100  2245.91 7.50  448.07 -2.26
 109 50 245 2093   836   0.97 100  1721.24 5.50  384.89 0.12
 110 50 245 2093   $*{}^{*}$   0.89 100  319.20 2.86  30.98 0.00
 111 50 367 672   $*{}^{*}$   0.57 0.00  0.49 0.35  2.62 0.00
 112 50 367 672   $*{}^{*}$   0.63 0.00  4.10 0.00  4.66 0.00
 113 50 367 672   $*{}^{*}$   0.68 0.00  0.53 0.00  2.43 0.00
 114 50 367 672   $*{}^{*}$   0.56 0.00  0.09 0.00  0.09 0.00
 115 50 367 672   $*{}^{*}$   0.64 0.36  0.44 0.18  2.59 0.00
 116 50 367 2687   $*{}^{*}$   1.52 7.35  34.96 0.33  50.68 0.00
 117 50 367 2687   $*{}^{*}$   1.44 7.80  34.21 0.81  28.74 0.00
 118 50 367 2687   $*{}^{*}$   1.50 8.18  23.60 1.53  5.18 0.34
 119 50 367 2687   $*{}^{*}$   1.54 13.72  34.00 1.74  15.30 0.00
 120 50 367 2687   $*{}^{*}$   1.57 7.00  15.72 2.33  8.92 0.00
 121 50 367 4702   726   2.26 100  3400.73 13.36  644.79 3.17
 122 50 367 4702   770   2.19 100  2418.38 16.36  870.81 0.52
 123 50 367 4702   786   2.30 100  2982.65 7.25  861.61 -0.64
 124 50 367 4702   711   2.42 100  952.21 4.92  929.33 0.28
 125 50 367 4702   764   2.16 13.61  1002.78 8.90  909.84 3.80
 126 50 490 1199   $*{}^{*}$   1.30 0.73  29.42 0.00  31.81 0.00
 127 50 490 1199   $*{}^{*}$   1.21 0.19  214.98 0.00  130.96 0.00
 128 50 490 1199   $*{}^{*}$   1.13 0.00  0.12 0.00  0.10 0.00
 129 50 490 1199   $*{}^{*}$   1.29 0.19  0.10 0.00  0.08 0.00
 130 50 490 1199   $*{}^{*}$   1.10 0.00  56.15 0.00  477.55 0.19
 131 50 490 4793   $*{}^{*}$   2.98 5.89  40.07 0.34  21.67 0.00
 132 50 490 4793   $*{}^{*}$   3.15 12.26  69.87 0.00  98.57 0.00
 133 50 490 4793   $*{}^{*}$   3.05 11.54  17.78 0.17  12.10 0.00
 134 50 490 4793   $*{}^{*}$   3.05 11.44  29.16 2.25  7.48 0.00
 135 50 490 4793   $*{}^{*}$   3.03 13.18  36.84 0.51  24.37 0.00
 136 50 490 8387   678   4.20 19.76  2445.19 16.96  449.03 -2.95
 137 50 490 8387   651   4.13 100  3606.37 15.36  964.71 0.61
 138 50 490 8387   689   4.27 100  1893.81 1.16  515.88 0.15
 139 50 490 8387   682   4.57 100  1301.15 9.09  749.00 2.05
 140 50 490 8387   674   4.47 22.85  2270.21 6.38  569.03 3.71
 141 75 555 1538   $*{}^{*}$   1.96 0.12  25.20 0.23  17.32 0.12
 142 75 555 1538   $*{}^{*}$   2.28 0.80  3.38 0.00  9.34 0.00
 143 75 555 1538   $*{}^{*}$   2.33 0.72  0.84 0.00  7.01 0.00
 144 75 555 1538   $*{}^{*}$   1.79 0.00  18.13 0.12  13.69 0.12
 145 75 555 1538   $*{}^{*}$   2.01 0.23  0.96 0.23  7.25 0.00
 146 75 555 6150   1047   5.64 18.05  2511.64 6.21  913.09 1.72
 147 75 555 6150   1069   5.57 12.91  2310.07 6.08  1038.47 -0.37
 148 75 555 6150   1040   5.50 100  3805.55 20.87  888.57 1.54
 149 75 555 6150   998   5.05 100  1086.86 1.10  543.73 -0.30
 150 75 555 6150   994   5.69 100  2127.32 18.01  1074.30 2.11
 156 75 832 3457   $*{}^{*}$   5.10 0.63  1.45 0.00  14.92 0.00
 157 75 832 3457   $*{}^{*}$   5.87 1.34  1.32 0.00  149.17 0.00
 158 75 832 3457   $*{}^{*}$   5.66 0.49  1.54 0.00  13.70 0.00
 159 75 832 3457   $*{}^{*}$   5.12 0.24  81.59 0.12  22.03 0.12
 160 75 832 3457   $*{}^{*}$   5.91 1.47  68.23 0.12  286.11 0.12
 161 75 832 13828   903   14.42 25.25  1977.61 -0.55  930.61 -0.33
 162 75 832 13828   953   13.11 18.05  3716.49 0.63  639.04 -1.57
 163 75 832 13828   892   13.52 16.03  3795.01 2.02  649.49 0.90
 164 75 832 13828   915   14.10 24.59  3343.98 1.75  1044.55 2.08
 165 75 832 13828   896   13.90 17.97  2741.60 4.24  856.37 0.89
 171 75 1110 6155   $*{}^{*}$   12.41 0.13  9.65 0.00  23.50 0.00
 172 75 1110 6155   $*{}^{*}$   10.83 0.13  50.48 0.00  51.93 0.00
 173 75 1110 6155   $*{}^{*}$   12.67 2.17  2.18 0.00  65.67 0.00
 174 75 1110 6155   $*{}^{*}$   12.30 0.64  3.04 0.13  28.39 0.00
 175 75 1110 6155   $*{}^{*}$   11.92 1.51  3.41 0.00  85.96 0.00
 176 75 1110 24620   867   27.07 19.49  2841.94 -0.35  755.04 -0.92
 177 75 1110 24620   851   29.82 23.15  3817.10 3.64  799.11 -0.12
 178 75 1110 24620   892   30.08 23.09  2969.99 6.84  1559.67 -0.90
 179 75 1110 24620   864   30.47 24.54  3356.33 2.20  895.40 0.81
 180 75 1110 24620   882   29.08 21.66  3096.31 -0.45  1572.95 -1.02
 183 75 1110 43085   1194   34.48 100  3619.07 100  3132.36 100
 186 100 990 4896   $*{}^{*}$   11.52 3.93  141.09 0.00  45.91 0.00
 187 100 990 4896   $*{}^{*}$   10.44 1.67  168.84 0.00  30.45 0.00
 188 100 990 4896   $*{}^{*}$   10.59 2.70  138.40 0.00  200.09 0.00
 189 100 990 4896   $*{}^{*}$   11.09 4.05  104.98 0.09  38.53 0.09
 190 100 990 4896   $*{}^{*}$   10.54 2.20  11.84 0.09  37.97 0.09
 191 100 990 19583   -   23.31 100  3847.28 100  1207.81 -100
 192 100 990 19583   1491   24.13 100  3832.03 100  2404.22 100
 193 100 990 19583   1510   23.93 100  3850.73 100  2060.18 100
 194 100 990 19583   1441   23.94 100  3980.25 100  1965.16 3.40
 195 100 990 19583   1560   24.58 100  3939.26 100  2594.71 5.71
 201 100 1485 11019   $*{}^{*}$   28.92 5.28  150.34 0.00  85.50 0.00
 202 100 1485 11019   $*{}^{*}$   32.30 2.65  107.84 0.00  218.25 0.00
 203 100 1485 11019   $*{}^{*}$   29.54 1.70  8.75 0.28  88.06 0.19
 204 100 1485 11019   $*{}^{*}$   30.32 1.24  115.85 1.05  142.18 0.29
 205 100 1485 11019   $*{}^{*}$   28.53 3.45  261.44 0.00  423.03 0.00
 206 100 1485 44075   1374   70.00 100  3846.93 100  3754.13 1.60
 207 100 1485 44075   1291   66.76 100  3848.22 100  3329.97 4.65
 208 100 1485 44075   1344   66.37 100  3836.87 100  3649.34 5.73
 209 100 1485 44075   1286   66.20 100  3858.32 100  3465.60 5.29
 210 100 1485 44075   1370   62.92 100  3844.58 100  3804.37 -2.26
 216 100 1980 19593   $*{}^{*}$   65.49 3.69  140.15 0.00  595.76 0.10
 217 100 1980 19593   $*{}^{*}$   70.46 4.34  95.81 0.00  118.17 0.10
 218 100 1980 19593   $*{}^{*}$   66.73 3.91  122.80 0.00  129.88 0.00
 219 100 1980 19593   $*{}^{*}$   61.12 1.76  82.52 0.00  113.50 0.00
 220 100 1980 19593   $*{}^{*}$   68.04 4.67  153.57 0.00  117.89 0.00
 221 100 1980 78369   1234   140.14 100  3893.96 100  3913.30 0.73
 222 100 1980 78369   1187   149.79 41.53  3697.46 100  3502.92 1.18
 223 100 1980 78369   1213   139.31 100  4020.74 100  3864.52 8.57
 224 100 1980 78369   1221   140.17 100  3999.16 100  2515.26 3.11
 225 100 1980 78369   $*{}^{*}$   127.94 35.42  3856.70 100  4061.45 2.25
  Average values   13.82 20.95  846.27 10.34  464.41 0.38
  %best   29.8   49.0   68.2
We notice that, for all the types of instances considered, KS returns the best average gap, as well as the highest percentage of instances with a non-positive gap with respect to the best solution value found in the literature. The 68.2% of the CCPR instances are solved with a non-positive gap, with an average gap of 0.38%, and an average computing time of less than $8$ minutes. It is definitely higher than the average computing time of the HDA+ method, which, however, returns an average gap of 20.95%. The classic KS approach is slower than KS, returning a gap of 10.34%, which is still less than the one returned by the HDA+ method.
It is clear from the comparison between the results of our KS approach and the classic KS one that using the maximum independent set problem, as well as the GRASP algorithm within the KS method, is crucial to address the MSTC problem. The increase in the computational time with respect to other existing methods is outweighed by the considerable reduction in the performance gap.

### v2

@@ 7 Conclusion
In this paper, we introduced a novel solution approach to the Minimum Spanning Tree with Conflicts problem, leveraging the KS method [1]. This algorithm focuses on obtaining a feasible solution, ideally of high quality, by strategically selecting a small set of promising variables. This subset, known as the kernel set, is initially constructed through the solution of the LP relaxation of the original ILP. Throughout the algorithm, the kernel set is dynamically enriched by identifying new promising variables through the solution of ILP subproblems of increasing size. A key innovation in our methodology lies in the incorporation of independent sets of proper subgraphs within the KS algorithm. This novel approach draws inspiration from the observation that the edges of any feasible MSTC solution correspond to an independent set in the conflict graph. The independent sets are used in order to both detect affinities among variables and find a starting point for the algorithm.
Through comparative analysis with existing heuristics, we demonstrated the efficacy of our method in producing competitive results. It is evident that our tailored KS approach offers a promising tool for tackling the MSTC problem and potentially other conflict-based optimization challenges.
Appendix A CCPR results
In this section, we show the detailed results of the comparison among the HDA+, ClassicKS, and KS on the CCPR instances proposed in [7]. As already noted for the other tables, whenever the B&C method finds the optimal value, we report such value with the “*” symbol in the UB_best column. Whenever no feasible solution is known in the literature, but we do find one, a “-” is reported in the UB_best column. In the gap columns, if the considered method does not find a feasible solution, we report a $100$. Instead, if the considered method does find a feasible solution, but there is a “-” in the UB_best column, we report a $-100$. Also in this case like for the previous tables, we report the results only on the instances where at least one of the methods (including ours) finds a feasible solution. As a consequence, the instances with ID 151–155, 166–170, 181, 182, 184, 185, 196–200, 211–215, 226–230 are not present in this table. The headings of Table LABEL:tab:kernel_cmst are the same as in Table 3. Again, we report the “Average values” and the “%best” rows at the end of the table.
@@ Appendix A CCPR results
   ID $n$ $m$ $|C|$     UB_best     time gap    time gap    time UB_time gap
   51 25 60 18     347     0.03 0.00    0.02 0.00    0.02 0.02 0.00
   52 25 60 18     389     0.03 0.00    0.00 0.00    0.00 0.00 0.00
   53 25 60 18     353     0.03 0.00    4.15 1.98    3.16 0.95 0.00
   54 25 60 18     346     0.03 0.00    0.00 0.00    0.00 0.00 0.00
   55 25 60 18     336     0.03 0.00    0.00 0.00    0.00 0.00 0.00
   56 25 60 71     381     0.04 0.00    0.42 0.00    0.15 0.11 0.00
   57 25 60 71     390     0.04 0.00    0.62 0.26    0.77 0.09 0.26
   58 25 60 71     372     0.04 0.00    0.02 0.00    0.04 0.04 0.00
   59 25 60 71     357     0.04 0.00    0.02 0.00    0.04 0.04 0.00
   60 25 60 71     406     0.04 0.00    0.00 0.00    0.00 0.00 0.00
   61 25 60 124     385     0.04 0.00    0.00 0.00    0.00 0.00 0.00
   62 25 60 124     432     0.04 0.00    0.00 0.00    0.00 0.00 0.00
   63 25 60 124     458     0.06 3.49    0.20 0.00    0.32 0.21 0.00
   64 25 60 124     400     0.05 0.00    0.26 1.25    0.07 0.04 0.75
   65 25 60 124     420     0.08 0.24    0.12 0.00    0.12 0.08 0.00
   66 25 90 41     311     0.03 0.00    2.80 0.00    0.25 0.10 0.00
   67 25 90 41     306     0.04 0.00    0.00 0.00    0.00 0.00 0.00
   68 25 90 41     299     0.03 0.00    0.69 0.00    0.24 0.14 0.00
   69 25 90 41     297     0.03 0.00    5.08 4.38    0.75 0.47 0.00
   70 25 90 41     318     0.03 0.00    12.87 0.00    5.95 0.79 0.00
   71 25 90 161     305     0.06 0.00    0.05 0.00    0.05 0.04 0.00
   72 25 90 161     339     0.07 0.00    0.00 0.00    0.00 0.00 0.00
   73 25 90 161     344     0.06 0.00    0.02 0.00    0.04 0.06 0.00
   74 25 90 161     329     0.07 0.61    0.10 0.30    0.24 0.20 0.00
   75 25 90 161     326     0.07 0.31    0.59 1.53    2.74 1.26 0.00
   76 25 90 281     349     0.08 0.00    0.25 0.00    0.41 0.34 0.00
   77 25 90 281     385     0.10 0.00    0.91 0.00    0.22 0.14 0.00
   78 25 90 281     335     0.09 0.00    0.17 0.30    0.20 0.12 0.00
   79 25 90 281     348     0.12 2.87    0.22 0.00    0.21 0.11 0.00
   80 25 90 281     357     0.11 0.56    0.16 0.00    0.22 0.15 0.00
   81 25 120 72     282     0.04 0.00    0.00 0.00    0.00 0.00 0.00
   82 25 120 72     294     0.04 0.00    0.97 0.00    15.40 3.44 0.00
   83 25 120 72     284     0.05 0.00    0.00 0.00    0.00 0.00 0.00
   84 25 120 72     281     0.04 0.00    2.56 0.00    1.90 0.10 0.00
   85 25 120 72     292     0.04 0.00    0.00 0.00    0.00 0.00 0.00
   86 25 120 286     321     0.10 0.00    0.14 0.31    0.23 0.11 0.00
   87 25 120 286     317     0.11 0.00    0.51 0.00    0.21 0.08 0.00
   88 25 120 286     284     0.08 0.00    0.00 0.00    0.00 0.00 0.00
   89 25 120 286     311     0.11 0.32    0.07 0.00    0.03 0.03 0.00
   90 25 120 286     290     0.08 0.00    0.02 0.00    0.06 0.06 0.00
   91 25 120 500     329     0.17 3.65    0.71 1.82    0.24 0.19 0.00
   92 25 120 500     339     0.16 2.36    2.83 0.00    0.63 0.48 0.00
   93 25 120 500     368     0.19 4.08    1.50 0.82    0.31 0.18 0.00
   94 25 120 500     311     0.14 0.97    0.14 0.00    0.38 0.11 0.00
   95 25 120 500     321     0.16 1.25    0.19 0.62    0.57 0.12 0.62
   96 50 245 299     619     0.24 0.00    4.27 0.00    442.27 71.54 0.00
   97 50 245 299     604     0.24 0.00    0.29 0.00    57.32 13.16 0.00
   98 50 245 299     634     0.22 0.00    0.01 0.00    0.01 0.01 0.00
   99 50 245 299     616     0.26 0.00    0.35 0.16    6.72 2.64 0.00
   100 50 245 299     595     0.28 0.00    21.59 0.00    0.25 0.25 0.00
   101 50 245 1196     678     0.62 2.95    10.90 0.00    24.81 5.09 0.00
   102 50 245 1196     681     0.64 5.87    86.60 3.08    26.29 8.49 0.00
   103 50 245 1196     709     0.64 2.26    17.80 0.71    12.86 3.83 0.00
   104 50 245 1196     639     0.55 2.66    11.05 0.16    6.51 2.19 0.00
   105 50 245 1196     681     0.65 9.84    23.49 0.59    15.44 4.55 0.29
   106 50 245 2093     833     1.00 100.00    1102.05 1.20    149.16 9.95 -2.04
   107 50 245 2093     835     1.02 100.00    3685.50 100.00    201.32 10.71 0.00
   108 50 245 2093     840     0.91 100.00    2245.91 7.50    79.02 9.81 -3.57
   109 50 245 2093     836     0.97 100.00    1721.24 5.50    34.16 17.88 -0.48
   110 50 245 2093     769     0.89 100.00    319.20 2.86    50.74 16.93 0.00
   111 50 367 672     570     0.57 0.00    0.49 0.35    105.67 21.60 0.00
   112 50 367 672     561     0.63 0.00    4.10 0.00    750.19 77.86 0.00
   113 50 367 672     573     0.68 0.00    0.53 0.00    8.60 2.73 0.00
   114 50 367 672     560     0.56 0.00    0.09 0.00    0.88 0.66 0.00
   115 50 367 672     549     0.64 0.36    0.44 0.18    203.51 0.29 0.00
   116 50 367 2687     612     1.52 7.35    34.96 0.33    277.78 44.35 0.00
   117 50 367 2687     615     1.44 7.81    34.21 0.81    19.76 4.04 0.00
   118 50 367 2687     587     1.50 8.18    23.60 1.53    2.46 0.97 0.00
   119 50 367 2687     634     1.54 13.72    34.00 1.74    86.84 16.41 0.00
   120 50 367 2687     643     1.57 7.00    15.72 2.33    8.08 3.19 0.00
   121 50 367 4702     726     2.26 100.00    3400.73 13.36    632.18 72.62 1.10
   122 50 367 4702     770     2.19 100.00    2418.38 16.36    116.46 15.51 -1.30
   123 50 367 4702     786     2.30 100.00    2982.65 7.25    483.81 57.09 -2.54
   124 50 367 4702     711     2.42 100.00    952.21 4.92    280.06 58.20 -1.97
   125 50 367 4702     764     2.16 13.61    1002.78 8.90    360.74 27.39 1.70
   126 50 490 1199     548     1.30 0.73    29.42 0.00    118.80 14.62 0.00
   127 50 490 1199     530     1.21 0.19    214.98 0.00    883.20 50.26 0.00
   128 50 490 1199     549     1.13 0.00    0.12 0.00    181.96 19.68 0.00
   129 50 490 1199     540     1.29 0.19    0.10 0.00    0.09 0.08 0.00
   130 50 490 1199     540     1.10 0.00    56.15 0.00    120.56 0.51 0.19
   131 50 490 4793     594     2.98 5.89    40.07 0.34    131.50 26.89 0.00
   132 50 490 4793     579     3.15 12.26    69.87 0.00    23.29 3.00 0.00
   133 50 490 4793     589     3.05 11.55    17.78 0.17    4.69 2.80 0.00
   134 50 490 4793     577     3.05 11.44    29.16 2.25    4.88 2.87 0.00
   135 50 490 4793     592     3.03 13.18    36.84 0.51    71.11 13.86 0.00
   136 50 490 8387     678     4.20 19.76    2445.19 16.96    476.13 48.62 -1.18
   137 50 490 8387     651     4.13 100.00    3606.37 15.36    998.68 155.43 -0.61
   138 50 490 8387     689     4.27 100.00    1893.81 1.16    396.10 76.11 -1.45
   139 50 490 8387     682     4.57 100.00    1301.15 9.09    290.68 42.49 0.73
   140 50 490 8387     674     4.47 22.85    2270.21 6.38    936.19 175.34 -2.52
   141 75 555 1538     868     1.96 0.12    25.20 0.23    573.26 30.45 0.00
   142 75 555 1538     871     2.28 0.80    3.38 0.00    772.83 99.49 0.00
   143 75 555 1538     838     2.33 0.72    0.84 0.00    19.99 7.48 0.00
   144 75 555 1538     855     1.79 0.00    18.13 0.12    638.29 104.89 0.00
   145 75 555 1538     857     2.01 0.23    0.96 0.23    10.30 1.32 0.00
   146 75 555 6150     1047     5.64 18.05    2511.64 6.21    400.59 43.88 0.48
   147 75 555 6150     1069     5.57 12.91    2310.07 6.08    1033.18 216.67 -2.15
   148 75 555 6150     1040     5.50 100.00    3805.55 20.87    754.06 95.82 -0.58
   149 75 555 6150     998     5.05 100.00    1086.86 1.10    852.15 128.74 1.40
   150 75 555 6150     994     5.69 100.00    2127.32 18.01    1039.43 151.19 0.40
   156 75 832 3457     798     5.10 0.63    1.45 0.00    609.13 103.84 0.00
   157 75 832 3457     821     5.87 1.34    1.32 0.00    301.01 39.39 0.00
   158 75 832 3457     816     5.66 0.49    1.54 0.00    522.14 28.66 0.00
   159 75 832 3457     820     5.12 0.24    81.59 0.12    654.14 49.56 0.12
   160 75 832 3457     815     5.91 1.47    68.23 0.12    297.86 30.99 0.00
   161 75 832 13828     903     14.42 25.25    1977.61 -0.55    1071.69 211.42 0.44
   162 75 832 13828     953     13.11 18.05    3716.49 0.63    943.99 136.26 -2.52
   163 75 832 13828     892     13.52 16.03    3795.01 2.02    1221.56 260.01 0.56
   164 75 832 13828     915     14.10 24.59    3343.98 1.75    1013.60 153.48 0.98
   165 75 832 13828     896     13.90 17.97    2741.60 4.24    434.95 61.89 1.12
   171 75 1110 6155     787     12.41 0.13    9.65 0.00    4.66 2.63 0.00
   172 75 1110 6155     785     10.83 0.13    50.48 0.00    257.80 46.85 0.00
   173 75 1110 6155     783     12.67 2.17    2.18 0.00    1297.35 118.07 0.00
   174 75 1110 6155     784     12.30 0.64    3.04 0.13    391.55 43.63 0.00
   175 75 1110 6155     797     11.92 1.51    3.41 0.00    33.91 7.69 0.00
   176 75 1110 24620     867     27.07 19.49    2841.94 -0.35    1290.64 240.51 1.15
   177 75 1110 24620     851     29.82 23.15    3817.10 3.64    1318.79 85.64 2.82
   178 75 1110 24620     892     30.08 23.09    2969.99 6.84    1115.62 162.08 -1.23
   179 75 1110 24620     864     30.47 24.54    3356.33 2.20    1373.69 150.56 11.81
   180 75 1110 24620     882     29.08 21.66    3096.31 -0.45    756.48 32.41 3.74
   183 75 1110 43085     1194     34.48 100.00    3619.07 100.00    1699.45 47.18 100.00
   186 100 990 4896     1119     11.52 3.93    141.09 0.00    834.38 48.88 0.00
   187 100 990 4896     1137     10.44 1.67    168.84 0.00    408.35 63.67 0.00
   188 100 990 4896     1113     10.59 2.70    138.40 0.00    1537.76 266.18 0.09
   189 100 990 4896     1110     11.09 4.05    104.98 0.09    684.33 141.51 0.00
   190 100 990 4896     1090     10.54 2.20    11.84 0.09    1284.48 167.24 0.09
   191 100 990 19583     -     23.31 100.00    3847.28 100.00    1290.60 140.05 -100.00
   192 100 990 19583     1491     24.13 100.00    3832.03 100.00    1159.85 44.61 17.91
   193 100 990 19583     1510     23.93 100.00    3850.73 100.00    1151.45 43.71 14.97
   194 100 990 19583     1441     23.94 100.00    3980.25 100.00    1149.78 42.00 23.73
   195 100 990 19583     1560     24.58 100.00    3939.26 100.00    1344.19 174.61 9.49
   201 100 1485 11019     1079     28.92 5.28    150.34 0.00    1898.20 301.64 0.19
   202 100 1485 11019     1056     32.30 2.65    107.84 0.00    935.21 183.48 0.00
   203 100 1485 11019     1059     29.54 1.70    8.75 0.28    761.93 100.02 0.00
   204 100 1485 11019     1046     30.32 1.24    115.85 1.05    871.40 119.46 0.00
   205 100 1485 11019     1072     28.53 3.45    261.44 0.00    1258.39 43.68 0.00
   206 100 1485 44075     1374     70.00 100.00    3846.93 100.00    1252.91 42.30 10.04
   207 100 1485 44075     1291     66.76 100.00    3848.22 100.00    1173.60 30.59 21.84
   208 100 1485 44075     1344     66.37 100.00    3836.87 100.00    1277.02 34.77 24.55
   209 100 1485 44075     1286     66.20 100.00    3858.32 100.00    1234.78 45.74 20.37
   210 100 1485 44075     1370     62.92 100.00    3844.58 100.00    1162.74 30.23 14.60
   216 100 1980 19593     1031     65.49 3.69    140.15 0.00    1478.66 16.70 0.00
   217 100 1980 19593     1036     70.46 4.34    95.81 0.00    275.72 46.57 0.00
   218 100 1980 19593     1024     66.73 3.91    122.80 0.00    761.29 101.41 0.00
   219 100 1980 19593     1025     61.12 1.76    82.52 0.00    1795.81 294.19 0.00
   220 100 1980 19593     1028     68.04 4.67    153.57 0.00    1064.45 159.70 0.00
   221 100 1980 78369     1234     140.14 100.00    3893.96 100.00    1516.22 102.96 20.10
   222 100 1980 78369     1187     149.79 41.53    3697.46 100.00    1283.16 54.50 25.61
   223 100 1980 78369     1213     139.31 100.00    3999.99 100.00    1310.06 90.75 21.85
   224 100 1980 78369     1221     140.17 100.00    3999.16 100.00    1311.85 45.53 17.85
   225 100 1980 78369     1245     127.94 35.42    3856.70 100.00    1274.09 43.89 11.65
    Average values     14.16 23.05    924.04 12.72    440.83 47.26 1.73
    %best     29.80     49.01     74.83
Table A1: Comparison of the results on the CCPR instances proposed in [7], for which at least a feasible solution is found.
It is clear from the comparison between the results of our KS approach and the classic KS one that using the maximum independent set problem, as well as the heuristic algorithm, is crucial to address the MSTC problem. The increase in the computational time with respect to other existing methods is outweighed by the considerable reduction in the performance gap.